OpenStudy (anonymous):

the product of two numbers is 100. if one decreased by 5 and the other increased by 1, the product is still 100. how do i find the numbers?

ganeshie8 (ganeshie8):

Let the two numbers be \(x\) and \(y\)

ganeshie8 (ganeshie8):

Can you set up equations based on given info ?

OpenStudy (anonymous):

xy= 100 (x-5)(y+1)=100 I'm not sure where to go from there. I set them equal to each other, but that didn't work

ganeshie8 (ganeshie8):

Your equations look good. Try eliminating one variable by substitution

ganeshie8 (ganeshie8):

\(xy= 100 \implies y = 100/x\) plug this in second equation

ganeshie8 (ganeshie8):

\((x-5)(\frac{100}{x}+1)=100\)

ganeshie8 (ganeshie8):

\((x-5)(\frac{100+x}{x})=100\)

ganeshie8 (ganeshie8):

\((x-5)(100+x)=100x\)

ganeshie8 (ganeshie8):

expand the left hand side you will see a quadratic equation which can be solved using quadratic formula

OpenStudy (anonymous):

x=-20 or x= 25. I understand it now. Thanks! Does the equation also work if I substitute it after I expand the binomials?

OpenStudy (anonymous):

I see that it does work. I did that in the beginning but I made an error somewhere, so I couldn't solve it. Thanks again.

ganeshie8 (ganeshie8):

np :)